OpenStudy (anonymous):

Find the height of a cylinder with radius 5 cm and surface area 1250 cm2.

OpenStudy (anonymous):

can u help me out plss?

OpenStudy (anonymous):

the formula for surface area of cylinder is 2\[2\pi r^{2} + 2\pi rh\] you already have the radius, so just substitute it into the formula

OpenStudy (anonymous):

ignore the 2 in the 1st sentence

OpenStudy (anonymous):

but how do i do it ? im confused!

OpenStudy (anonymous):

let 2πr2+2πrh = 1250 then, h = (1250- 2π(5)^2)/ 2π(5)^2) where 5 is your radius

OpenStudy (anonymous):

oops h = (1250- 2π(5)^2)/ 2π(5))

OpenStudy (anonymous):

so is going to be 5 right ?

OpenStudy (anonymous):

si is going to be 13.42 right ?

OpenStudy (anonymous):

h = (1250 - (2 * 3.142 * 25 )) / (2 * 3.142 * 5) = 1250 - 157.14 / 31.42 = 1092.86 / 31.42 = 34.782